Taylor is 3 years older than Nova. In 5 years, twice the sum of their ages will equal 5 times Nova's age today. How old are Taylor and Nova today? show all work
Answer by Paul(988)

Let nova's age be represented by x and taylor's by 3+x
(3+x)+x
5+6+2x+2x=5x
5x=11+4x
x=11
11+4=15. hence Nova is 11 and Taylor is 15